U Brew blends Colombian coffee worth $0.93 per pound with Kenyan coffee worth $1.20 per pound to get 30 pounds of a coffee blend worth $1.02 per pound. How many pounds of each type of coffee are used?
Answer by richwmiller(17219)

look in my problems I solved this yesterday
x+y=30,
.93x+1.2y=1.02*30
x=20
y=10